### v2

# A Generalization of Zalcman’s Lemma on Complex Lie Groups

Abstract.
Zalcman’s Lemma makes significant applications in normal families, complex dynamics and related problems in complex analysis. In the present paper, we are devoted to generalizing the classical Zalcman’s lemma to complex Lie groups by means of exponential mappings defined by holomorphic one-parameter subgroups.
Key words and phrases:
Normal families; Zalcman’s lemma; Marty’s criterion; Heuristic principle; Complex Lie groups
2010 Mathematics Subject Classification:
32A19; 32H02; 32H25
Correspondence author
1. Introduction
In 1975, Zalcman [22] contributed an important result in characterization of normality of families of meromorphic functions on $\mathbb{C}.$ He proved that
Theorem A (Zalcman’s lemma).
Let $\mathscr{F}$ be a family of meromorphic functions on a domain $\Omega\subset\mathbb{C}.$ Then, $\mathscr{F}$ is not normal on $\Omega$ if and only if there exist
$(a)$ a sequence $\{f_{j}\}_{j=1}^{+\infty}\subset\mathscr{F};$
$(b)$ a sequence $\{p_{j}\}_{j=1}^{+\infty}\subset\Omega$ with $p_{j}\to p_{0}\in\Omega$ as $j\to+\infty;$
$(c)$ a sequence $\{\rho_{j}\}_{j=1}^{+\infty}\subset\mathbb{R}$ with $0<\rho_{j}\to 0$ as $j\to+\infty$
such that the sequence
$\phi_{j}(z):=f_{j}(p_{j}+\rho_{j}z),\ \ \ \ j=1,2,\cdots$
converges uniformly on compact subsets of $\mathbb{C}$ to a nonconstant meromorphic function $\phi$ on $\mathbb{C}.$
Zalcman’s work was inspired by the result of Lohwater and Pommerenke [10], which establishes a well-known heuristic principle showing that Bloch’s principle [2] is valid to a certain Picard type property. Due to the importance of Zalcman’s lemma, many authors have been working on its generalizations. A beautiful work on this lemma is due to Pang [13, 14] who proved a stronger version with a real parameter, which thus shows that Bloch’s principle holds for some Picard type properties related to derivatives. More details for this lemma and its applications in one complex variable refer to [4, 9, 16, 20, 23] and references therein.
Many efforts were made to generalize Zalcman’s lemma to several complex variables, and more generally to complex manifolds. An earlier study about this lemma was carried by Aladro and Krantz [1]. Unfortunately, their result (Theorem 3.1, [1]) was later pointed out to be incorrect by Thai, Trang and Huong [17] who presented a correct extension of Zalcman’s lemma as follows.
Theorem B.
Let $N$ be a compact Hermitian manifold, and $\Omega$ a domain in $\mathbb{C}^{m}.$ Let $\mathscr{F}:\Omega\to N$ be a family of holomorphic mappings. Then, $\mathscr{F}$ is not normal on $\Omega$ if and only if there exist
$(a)$ a sequence $\{f_{j}\}_{j=1}^{+\infty}\subset\mathscr{F};$
$(b)$ a sequence $\{p_{j}\}_{j=1}^{+\infty}\subset\Omega$ with $p_{j}\to p_{0}\in\Omega$ as $j\to+\infty;$
$(c)$ a sequence $\{\rho_{j}\}_{j=1}^{+\infty}\subset\mathbb{R}$ with $0<\rho_{j}\to 0$ as $j\to+\infty$
such that the sequence
$\phi_{j}(z):=f_{j}(p_{j}+\rho_{j}z),\ \ \ \ j=1,2,\cdots$
converges uniformly on compact subsets of $\mathbb{C}^{m}$ to a nonconstant holomorphic mapping $\phi:\mathbb{C}^{m}\to N.$
For $N=\mathbb{P}^{1}(\mathbb{C}),$ Theorem B was generalized by Chang, Xu and Yang [5], Dovbush [7], Yang and Pang [21], etc. More details refer to [6] and references therein.
Unfortunately, unlike Euclidean spaces, a topological manifold is generally not a linear space, which lacks addition and scalar multiplication operations. So, developing Zalcman’s lemma on a general complex manifold seems to be a non-trivial task. To our knowledge, there is no literature about Zalcman’s lemma that is established on a non-Euclidean complex manifold. To fill this gap, we are devoted to a generalization of this lemma from $\mathbb{C}^{m}$ to a complex Lie group of finite dimension. There are many classical examples of complex Lie groups such as $\mathbb{C}^{m},\mathbb{C}^{m}\setminus\Lambda,GL(m,\mathbb{C}),SL(m,\mathbb{C}),O(m,\mathbb{C}),$ etc., see [11].
Let $G$ be a finite-dimensional complex Lie group whose identity element is denoted by $e.$ Note that each $g\in G$ can define a left translation $L_{g}:G\to G$ by $L_{g}(h)=gh$ for $h\in G.$ Let $T^{1,0}_{g}G$ be the holomorphic tangent space of $G$ at $g\in G.$ The holomorphic Lie algebra of $G$ is defined by $\mathfrak{g}=T^{1,0}_{e}G.$ Let
$\exp_{g}:\ T^{1,0}_{g}G\to G,\ \ \ \ g\in G$
be the exponential mapping which is a holomorphic mapping defined by the holomorphic one-parameter subgroup of $G,$ see Section 2 for details.
We establish a version of Zalcman’s lemma on complex Lie groups.
Theorem 1.1.
Let $G$ be an $m$-dimensional complex Lie group with holomorphic Lie algebra $\mathfrak{g},$ and $\Omega$ a domain in $G.$ Fix a basis $\varepsilon=(\varepsilon_{1},\cdots,\varepsilon_{m})$ of $\mathfrak{g}.$ Let $N$ be a compact Hermitian manifold. Let $\mathscr{F}:\Omega\to N$ be a family of holomorphic mappings. Then, $\mathscr{F}$ is not normal on $\Omega$ if and only if there exist
$(a)$ a sequence $\{f_{j}\}_{j=1}^{+\infty}\subset\mathscr{F};$
$(b)$ a sequence $\{p_{j}\}_{j=1}^{+\infty}\subset\Omega$ with $p_{j}\to p_{0}\in\Omega$ as $j\to+\infty;$
$(c)$ a sequence $\{\rho_{j}\}_{j=1}^{+\infty}\subset\mathbb{R}$ with $0<\rho_{j}\to 0$ as $j\to+\infty$
such that the sequence
$\phi_{j}(z):=f_{j}\circ\exp_{p_{j}}\big(\rho_{j}(dL_{p_{j}})_{e}(\varepsilon z^{T})\big),\ \ \ \ j=1,2,\cdots$
converges uniformly on compact subsets of $\mathbb{C}^{m}$ to a nonconstant holomorphic mapping $\phi:\mathbb{C}^{m}\to N.$ Here, $z^{T}$ is the transpose of $z=(z_{1},\cdots,z_{m}).$
In further, we obtain the following local version of Theorem 1.1.
Corollary 1.2.
Let $G$ be an $m$-dimensional complex Lie group with holomorphic Lie algebra $\mathfrak{g},$ and $\Omega$ a domain in $G.$ Fix a basis $\varepsilon=(\varepsilon_{1},\cdots,\varepsilon_{m})$ of $\mathfrak{g}.$ Let $N$ be a compact Hermitian manifold. Let $\mathscr{F}:\Omega\to N$ be a family of holomorphic mappings. Then, $\mathscr{F}$ is not normal at $p_{0}\in\Omega$ if and only if there exist
$(a)$ a sequence $\{f_{j}\}_{j=1}^{+\infty}\subset\mathscr{F};$
$(b)$ a sequence $\{p_{j}\}_{j=1}^{+\infty}\subset\Omega$ with $p_{j}\to p_{0}$ as $j\to+\infty;$
$(c)$ a sequence $\{\rho_{j}\}_{j=1}^{+\infty}\subset\mathbb{R}$ with $0<\rho_{j}\to 0$ as $j\to+\infty$
such that the sequence
$\phi_{j}(z):=f_{j}\circ\exp_{p_{j}}\big(\rho_{j}(dL_{p_{j}})_{e}(\varepsilon z^{T})\big),\ \ \ \ j=1,2,\cdots$
converges uniformly on compact subsets of $\mathbb{C}^{m}$ to a nonconstant holomorphic mapping $\phi:\mathbb{C}^{m}\to N.$ Here, $z^{T}$ is the transpose of $z=(z_{1},\cdots,z_{m}).$
To see how Theorem B can be covered by our result, we consider a special case where $G=\mathbb{C}^{m}.$ We regard $\mathbb{C}^{m}$ as an additive group with zero element $\textbf{0},$ and define the left translation $L_{p}(z)=p+z$ for $p,z\in\mathbb{C}^{m}.$ Put
$\varepsilon=\Big(\frac{\partial}{\partial\zeta_{1}},\cdots,\frac{\partial}{\partial\zeta_{m}}\Big)\Big|_{\textbf{0}},$
where $\zeta_{1},\cdots,\zeta_{m}$ stand for complex coordinates of $\mathbb{C}^{m}.$ Since $T^{1,0}_{z}\mathbb{C}^{m}\cong\mathbb{C}^{m}$ for all $z\in\mathbb{C}^{m},$ it is immediate that
$(dL_{p_{j}})_{\textbf{0}}(\varepsilon z^{T})=\Big(z_{1}\frac{\partial}{\partial\zeta_{1}}+\cdots+z_{m}\frac{\partial}{\partial\zeta_{m}}\Big)\Big|_{p_{j}}\in T^{1,0}_{p_{j}}\mathbb{C}^{m}.$
By the definition of $\exp,$ we obtain $\exp_{p_{j}}(\rho_{j}(dL_{p_{j}})_{\textbf{0}}(\varepsilon z^{T}))=p_{j}+\rho_{j}z,$ which indicates that Theorem 1.1 extends Theorem B.
Afterwards, we shall state a Robinson-Zalcman type heuristic principle of holomorphic families of a Picard type property on a complex Lie group. Let $g$ be a holomorphic mapping on a domain $D$ in a complex manifold. Denote by $P$ a property. Write $(g,D)\in P,$ if $g$ is of $P$ on $D;$ $(g,x)\in\neg P,$ if $g$ is not of $P$ in any neighborhood of $x$ in $D;$ and $(g,D)\in\neg P,$ if $(g,x)\in\neg P$ for all $x\in D.$
Let $\Omega$ be a domain in a complex Lie group $G,$ and $N$ a compact Hermitian manifold. Let $f,f_{j}:\Omega\to N$ be holomorphic mappings. For $p,p_{j}\in G,$ put
$\psi(z)=\exp_{p}(\theta z^{T});\ \ \ \ \ \phi_{j}(z)=f_{j}\circ\exp_{p_{j}}(\theta^{j}z^{T}),\ \ \ \ j=1,2,\cdots$
where $\theta,\theta^{j}$ are basis of $T^{1,0}_{p}G,T^{1,0}_{p_{j}}G,$ respectively. With these notations, we obtain a Robinson-Zalcman type heuristic principle as follows.
Theorem 1.3.
Let $G$ be an $m$-dimensional complex Lie group, and $\Delta,\Omega$ domains in $G$ with $\Delta\subset\Omega.$ Let $N$ be a compact Hermitian manifold. Assume that $P$ is a property of holomorphic mappings into $N$ such that
$(a)$ if $(f,\Omega)\in P,$ then $(f,\Delta)\in P$ for any $\Delta;$
$(b)$ if $(f,\Delta)\in P,$ then $(f\circ\psi,\psi^{-1}(\Delta))\in P$ for any $p,\theta;$
$(c)$ if $\{\phi_{j};(\phi_{j},D_{j})\in P\}_{j=1}^{+\infty}$ converges uniformly on compact subsets of $\mathbb{C}^{m}$ to a holomorphic mapping $\phi:\mathbb{C}^{m}\to N$ for domains $D_{1},D_{2},\cdots$ in $\mathbb{C}^{m}$ such that $D_{1}\subset D_{2}\subset\cdots,$ then $(\phi,\mathbb{C}^{m})\in P$ or $(\phi,\mathbb{C}^{m})\in\neg P;$
$(d)$ if $(\phi,\mathbb{C}^{m})\in P$ or $(\phi,\mathbb{C}^{m})\in\neg P,$ then $\phi$ is constant.
Then, the family $\mathscr{F}=\{f:(f,\Omega)\in P\}$ is normal on $\Omega.$
2. Exponential Mappings of Complex Lie Groups
A complex Lie group $G$ is a complex manifold with a group structure such that its multiplication $(g,h)\mapsto gh$ and inverse $g\mapsto g^{-1}$ are holomorphic for $g,h\in G.$ In this section, we shall prove some basic properties of exponential mappings of complex Lie groups. We refer the reader to [3, 8, 11, 15, 18] for a more in-depth understanding of real and complex Lie groups.
Let $G$ be a finite-dimensional complex Lie group with identity element $e.$ We use $\mathfrak{g}$ to denote the holomorphic Lie algebra of $G$ defined by $\mathfrak{g}:=T^{1,0}_{e}G,$ where $T^{1,0}_{e}G$ denotes the holomorphic tangent space of $G$ at $e.$ A holomorphic vector field $\tilde{X}$ on $G$ is called left-invariant, if $(dL_{g})_{h}\tilde{X}_{h}=\tilde{X}_{gh}$ for $g,h\in G,$ where $(dL_{g})_{h}$ is the differential or derivative of left translation $L_{g}$ at $h.$ Note that each vector $X\in\mathfrak{g}$ generates a unique left-invariant holomorphic vector field $\tilde{X}$ on $G$:
$\tilde{X}_{g}=(dL_{g})_{e}(X)\in T^{1,0}_{g}G,\ \ \ \ g\in G$
with $\tilde{X}_{e}=X.$ Associate a $\mathscr{C}^{1}$-continuous curve $\gamma:\mathbb{C}\to G$ satisfying
$\gamma^{\prime}=\tilde{X}_{\gamma},\ \ \ \ \gamma(0)=e.$
It is known that such a curve $\gamma$ has to be existent uniquely and holomorphic due to ODE theory. For convenience, we shall use $\gamma_{X}$ to denote this solution associated to $X.$
The exponential mapping $\exp:\mathfrak{g}\to G$ is defined by $\exp(X)=\gamma_{X}(1)$ with $X\in\mathfrak{g},$ where $\gamma_{X}:\mathbb{C}\to G$ is a holomorphic curve satisfying $\gamma^{\prime}_{X}=\tilde{X}_{\gamma_{X}}$ with $\gamma_{X}(0)=e.$ Clearly, $\exp$ is holomorphic on $\mathfrak{g}.$
Property 2.1.
$\{\gamma_{X}(z)\}_{z\in\mathbb{C}}$ is a holomorphic one-parameter subgroup.
Proof.
Since $\gamma_{X}(0)=e\in\mathfrak{g},$ it suffices to prove that $\gamma_{X}(\zeta)\gamma_{X}(z)=\gamma_{X}(\zeta+z)$ for $\zeta,z\in\mathbb{C}.$ Set $\alpha_{\zeta}(z)=\gamma_{X}(\zeta)\gamma_{X}(z)$ and $\beta_{\zeta}(z)=\gamma_{X}(\zeta+z).$ It is equivalent to show that $\alpha_{\zeta},\beta_{\zeta}$ solve the following Cauchy initial valued problem
(1) $\gamma^{\prime}=\tilde{X}_{\gamma},\ \ \ \ \gamma(0)=\gamma_{X}(\zeta)$
due to the uniqueness theorem. By $\gamma_{X}(\zeta)\gamma_{X}(z)=L_{\gamma_{X}(\zeta)}(\gamma_{X}(z)),$ we obtain
$\alpha^{\prime}_{\zeta}(z)$ $=$ $(dL_{\gamma_{X}(\zeta)})_{\gamma_{X}(z)}\gamma^{\prime}_{X}(z)$
$=$ $(dL_{\gamma_{X}(\zeta)})_{\gamma_{X}(z)}\tilde{X}_{\gamma_{X}(z)}$
$=$ $(dL_{\gamma_{X}(\zeta)})_{\gamma_{X}(z)}(dL_{\gamma_{X}(z)})_{e}X$
$=$ $(dL_{\gamma_{X}(\zeta)\gamma_{X}(z)})_{e}X$
$=$ $(dL_{\alpha_{\zeta}(z)})_{e}X$
$=$ $\tilde{X}_{\alpha_{\zeta}(z)}.$
Moreover, we have $\alpha_{\zeta}(0)=\gamma_{X}(\zeta)\gamma_{X}(0)=\gamma_{X}(\zeta)e=\gamma_{X}(\zeta).$ Hence, $\alpha_{\zeta}$ solves Eq. (1). Also, by $\beta_{\zeta}(0)=\gamma_{X}(\zeta)$ and $\beta^{\prime}_{\zeta}(z)=\gamma^{\prime}_{X}(\zeta+z)=\tilde{X}_{\gamma_{X}(\zeta+z)}=\tilde{X}_{\beta_{\zeta}(z)},$ we deduce that $\beta_{\zeta}$ solves Eq. (1). Hence, we have $\alpha_{\zeta}=\beta_{\zeta}.$
Property 2.2.
For any $X\in\mathfrak{g},$ we have
$\gamma_{X}(z)=\exp(zX),\ \ \ \ \ z\in\mathbb{C}.$
Proof.
Set $\alpha(t)=\gamma_{X}(zt),\beta(t)=\gamma_{zX}(t)$ for $t\in\mathbb{C}.$ Since $\exp(zX)=\gamma_{zX}(1),$ it is sufficient to prove that $\alpha(t)=\beta(t)$ by putting $t=1,$ which is equivalent to show that $\alpha,\beta$ solve the same Cauchy initial value problem. By
$\alpha^{\prime}(t)=z\gamma^{\prime}_{X}(zt)=z\tilde{X}_{\alpha(t)}=(dL_{\alpha(t)})_{e}(zX),$
$\beta^{\prime}(t)=\widetilde{zX}_{\beta(t)}=(dL_{\beta(t)})_{e}(zX),$
we deduce that $\alpha,\beta$ satisfy the same first-order differential equation. Since $\alpha(0)=\beta(0)=e,$ we conclude that $\alpha=\beta.$ This completes the proof. ∎
Property 2.3.
For any $X\in\mathfrak{g},$ we have
$\frac{d}{dz}\exp(zX)=\big(dL_{\exp(zX)}\big)_{e}(X),\ \ \ \ z\in\mathbb{C}.$
Proof.
By Property 2.2, we have $\exp(zX)=\gamma_{X}(z).$ Then
$\frac{d}{dz}\exp(zX)=\gamma^{\prime}_{X}(z)=\tilde{X}_{\gamma_{X}(z)}=\big(dL_{\gamma_{X}(z)}\big)_{e}(X)=\big(dL_{\exp(zX)}\big)_{e}(X).$
In general, a vector $\xi\in T_{g}^{1,0}G$ with $g\in G$ generates a unique left-invariant holomorphic vector field $\tilde{\xi}$ on $G$:
$\tilde{\xi}_{h}=(dL_{hg^{-1}})_{g}(\xi)=(dL_{h})_{e}\circ(dL_{g^{-1}})_{g}(\xi)\in T^{1,0}_{h}G,\ \ \ \ h\in G$
with $\tilde{\xi}_{g}=\xi.$ Similarly, we define the exponential mapping
$\exp_{g}:\ T_{g}^{1,0}G\to G,\ \ \ \ g\in G$
via $\exp_{g}(\xi)=\gamma_{\xi}(1)$ for $\xi\in T^{1,0}_{g}G,$ where $\gamma_{\xi}:\mathbb{C}\to G$ is a holomorphic curve solving uniquely the differential equation $\gamma^{\prime}=\tilde{\xi}_{\gamma}$ with initial value $\gamma(0)=g.$ It is clear that $\exp_{g}$ is holomorphic on $T_{g}^{1,0}G.$ For convenience, we shall use $\gamma_{\xi}$ to denote this solution associated to $\xi.$
Applying the similar arguments as in the proofs of Properties 2.1 and 2.2, we can also show the following Properties 2.4 and 2.5.
Property 2.4.
$\{\gamma_{\xi}(z)\}_{z\in\mathbb{C}}$ is a holomorphic one-parameter subgroup.
Property 2.5.
For any $\xi\in T_{g}^{1,0}G$ with $g\in G,$ we have
$\gamma_{\xi}(z)=\exp_{g}(z\xi),\ \ \ \ z\in\mathbb{C}.$
Property 2.6.
For any $\xi\in T^{1,0}_{g}G$ with $g\in G,$ we have
$\exp_{g}(z\xi)=g\exp(zX),\ \ \ \ X=(dL_{g^{-1}})_{g}(\xi).$
Proof.
Set $\eta(z)=g\exp(zX)=L_{g}\circ\exp(zX).$ By virtue of Property 2.5, we have $\exp_{g}(z\xi)=\gamma_{\xi}(z).$ Since $\eta(0)=\gamma_{\xi}(0)=g,$ it suffices to prove that $\eta,\gamma_{\xi}$ satisfy the same first-order differential equation. By Property 2.3
$\eta^{\prime}(z)$ $=$ $(dL_{g})_{\exp(zX)}\circ(dL_{\exp(zX)})_{e}(X)$
$=$ $d\big(L_{g}\circ L_{\exp(zX)}\big)_{e}(X)$
$=$ $\big(dL_{g\exp(zX)}\big)_{e}(X)$
$=$ $\big(dL_{\eta(z)}\big)_{e}(X).$
On the other hand, we have
$\gamma^{\prime}_{\xi}(z)=\tilde{\xi}_{\gamma_{\xi}(z)}=(dL_{\gamma_{\xi}(z)})_{e}\circ(dL_{g^{-1}})_{g}(\xi)=(dL_{\gamma_{\xi}(z)})_{e}(X).$
Therefore, $\eta,\gamma_{\xi}$ satisfy the same first-order differential equation. The proof is completed. ∎
If $z=1$ is taken in Property 2.6, then it is immediate that
Corollary 2.7.
For any $g\in G,$ we have
$\exp_{g}=L_{g}\circ\exp\circ(dL_{g^{-1}})_{g}.$
Fix $\xi\in T^{1,0}_{g}G$ with $g\in G.$ Let
$\iota_{\xi}:\ T^{1,0}_{g}G\to T^{1,0}_{\xi}(T^{1,0}_{g}G)$
be the canonical linear isomorphism, i.e., $\iota_{\xi}(\eta)=\theta^{\prime}_{\eta}(0)$ for $\eta\in T^{1,0}_{g}G,$ where $\theta_{\eta}(z)=\xi+z\eta$ is a holomorphic curve in $T^{1,0}_{g}G.$ For multiple tangent bundles of $G,$ one regards them as identical. Equip $G$ with a left-invariant Hermitian metric (such a metric is aways existent, see [8, 18]), i.e.,
$\langle\xi,\eta\rangle=\big\langle(dL_{h})_{g}(\xi),(dL_{h})_{g}(\eta)\big\rangle,\ \ \ \ \xi,\eta\in T^{1,0}_{g}G$
for any $g,h\in G.$ With this metric, $(dL_{h})_{g}$ is a linear isometry and so
(2) $\|(dL_{h})_{g}\|=\|(dL_{h})^{-1}_{g}\|=1,\ \ \ \ g,h\in G.$
This metric on $T^{1,0}G$ induces a natural metric on $T^{1,0}(T^{1,0}G)$ defined by
$\langle\zeta,\eta\rangle:=\big\langle\iota_{\xi}^{-1}(\zeta),\iota_{\xi}^{-1}(\eta)\big\rangle,\ \ \ \ \zeta,\eta\in T^{1,0}_{\xi}(T^{1,0}_{g}G),$
which forces $\iota_{\xi}$ to be a linear isometry and it is therefore
(3) $\|\iota_{\xi}\|=\|\iota_{\xi}^{-1}\|=1,\ \ \ \ \|d\iota_{\xi}\|=\|d\iota^{-1}_{\xi}\|=1.$
By (2) and (3), we may regard in order to simplify the formula that
(4) $(d(dL_{h})_{g})_{\xi}=(dL_{h})_{g}\circ\iota^{-1}_{\xi},$
up to a linear isometry.
Let
$(d\exp_{g})_{\xi}:\ T^{1,0}_{\xi}(T^{1,0}_{g}G)\to T^{1,0}_{\exp_{g}(\xi)}G$
be the differential of $\exp_{g}:T^{1,0}_{g}G\to G$ at $\xi,$ which can be represented as
$(d\exp_{g})_{\xi}(\eta)=\frac{d}{dz}\Big|_{z=0}\exp_{g}\big(\xi+z\iota_{\xi}^{-1}(\eta)\big),\ \ \ \ \eta\in T^{1,0}_{\xi}(T^{1,0}_{g}G).$
For $d\exp_{X},$ the well-known Duhamel formula (see, e.g., Theorem 5 in Section 1.2, [15]) gives an explicit expression
(5) $d\exp_{X}=\big(dL_{\exp(X)}\big)_{e}\circ\frac{Id-e^{-{\rm{ad}}_{X}}}{{\rm{ad}}_{X}}\circ\iota^{-1}_{X},\ \ \ \ X\in\mathfrak{g},$
where
$\frac{Id-e^{-{\rm{ad}}_{X}}}{{\rm{ad}}_{X}}=\sum_{k=0}^{+\infty}\frac{(-1)^{k}}{(k+1)!}({\rm{ad}}_{X})^{k}$
with ${\rm{ad}}_{X}(Y)=[X,Y]=XY-YX$ for $Y\in\mathfrak{g}.$ Set $X=(dL_{g^{-1}})_{g}(\xi).$ Using Corollary 2.7, (4) and (5), we derive that
$(d\exp_{g})_{\xi}$ $=$ $(dL_{g})_{\exp(X)}\circ\big(dL_{\exp(X)}\big)_{e}\circ\frac{Id-e^{-{\rm{ad}}_{X}}}{{\rm{ad}}_{X}}$
$\circ\iota^{-1}_{X}\circ(dL_{g^{-1}})_{g}\circ\iota^{-1}_{\xi}.$
Property 2.8.
For any $g\in G,$ $(d\exp_{g})_{\emph{{0}}}$ is a linear isometry.
Proof.
Using (2) (or combining Corollary 2.7 and (4) with (5)), we obtain
$(d\exp_{g})_{\textbf{0}}=(dL_{g})_{e}\circ\iota^{-1}_{\textbf{0}}\circ(dL_{g^{-1}})_{g}\circ\iota^{-1}_{\textbf{0}},$
which is a linear isometry. ∎
Define
$C_{\mathfrak{g}}:=\sup_{X,Y\in\mathfrak{g},\ \|X\|=\|Y\|=1}\|[X,Y]\|,$
which is called the structure constant of $G.$ Note that $0\leq C_{\mathfrak{g}}<+\infty,$ since $\mathfrak{g}$ has finite dimension. By $\|{\rm{ad}}_{X}(Y)\|=\|[X,Y]\|\leq C_{\mathfrak{g}}\|X\|\|Y\|,$ we get
(7) $\|{\rm{ad}}_{X}\|\leq C_{\mathfrak{g}}\|X\|.$
Theorem 2.9.
For any $\xi\in T^{1,0}_{g}G$ with $g\in G,$ we have
$\big\|(d\exp_{g})_{\xi}\big\|\leq\frac{e^{C_{\mathfrak{g}}\|\xi\|}-1}{C_{\mathfrak{g}}\|\xi\|}.$
Proof.
It yields from (2), (3), (2) and (7) that
$\big\|(d\exp_{g})_{\xi}\big\|$ $\leq$ $\big\|(dL_{g})_{\exp(X)}\big\|\cdot\big\|\big(dL_{\exp(X)}\big)_{e}\big\|\cdot\bigg\|\frac{Id-e^{-{\rm{ad}}_{X}}}{{\rm{ad}}_{X}}\bigg\|$
$\cdot\big\|\iota^{-1}_{X}\big\|\cdot\big\|(dL_{g^{-1}})_{g}\big\|\cdot\big\|\iota^{-1}_{\xi}\big\|$
$\leq$ $\bigg\|\frac{Id-e^{-{\rm{ad}}_{X}}}{{\rm{ad}}_{X}}\bigg\|$
$\leq$ $\sum_{k=0}^{+\infty}\frac{1}{(k+1)!}\big(C_{\mathfrak{g}}\|X\|\big)^{k}$
$=$ $\frac{e^{C_{\mathfrak{g}}\|\xi\|}-1}{C_{\mathfrak{g}}\|\xi\|},$
where $X=(dL_{g^{-1}})_{g}(\xi).$
3. Proofs and Picard Theorem
Let $\mathscr{C}(M,N)$ denote the set of all continuous mappings between complex manifolds $M,N.$ Equip $\mathscr{C}(M,N)$ with the compact-open topology. A family $\mathscr{F}\subset\mathscr{C}(M,N)$ is called normal, if and only if each sequence of $\mathscr{F}$ contains a subsequence which is either relatively compact in $\mathscr{C}(M,N)$ or compactly divergent (Definition 1.1, [19]). Let $\mathscr{H}(M,N)\subset\mathscr{C}(M,N)$ denote the set of all holomorphic mappings between $M,N.$ Anyway, once a distance function is chosen on $N,$ a sequence $\{f_{j}\}_{j=1}^{+\infty}\subset\mathscr{C}(M,N)$ converges to an $f\in\mathscr{C}(M,N)$ if and only if $\{f_{j}\}_{j=1}^{+\infty}$ converges uniformly on compact subsets to $f.$ Hence, a basic fact derived from Weierstrass theorem shows that $\mathscr{H}(M,N)$ is closed in $\mathscr{C}(M,N).$ When $N$ is a compact Hermitian manifold, using Ascoli-Arzelà theorem or a result of Wu (Lemma 1.1, [19]), $\mathscr{F}$ is normal on $M$ as long as $\mathscr{F}$ is equicontinuous on $M.$
To prove Theorem 1.1, we extend Marty’s criterion [12] for normality. Let $f:M\to N$ be a holomorphic mapping between Hermitian manifolds $M,N.$ The operator norm $\|df\|$ of differential $df$ is defined by
$\|df_{x}\|=\sup_{\xi\in T^{1,0}_{x}M}\frac{\|df_{x}(\xi)\|_{N}}{\|\xi\|_{M}},\ \ \ \ x\in M.$
We provide a slight extension of Marty’s criterion as follows.
Lemma 3.1.
Let $M,N$ be Hermitian manifolds with that $N$ is compact. Let $\mathscr{F}:M\to N$ be a family of holomorphic mappings. Then, $\mathscr{F}$ is normal on $M$ if and only if $\{\|df\|\}_{f\in\mathscr{F}}$ is locally uniformly bounded on $M.$
Proof.
We first prove the sufficiency. Let $x_{0}\in M$ be an arbitrary point, and take a small neighborhood $U(x_{0})$ of $x_{0}$ in $M.$ For a fixed $x\in U(x_{0}),$ we treat a $\mathscr{C}^{1}$-continuous curve $\gamma:[0,1]\to M$ which connects $x_{0}$ with $x.$ It is evident that $f\circ\gamma:[0,1]\to N$ is a $\mathscr{C}^{1}$-continuous curve connecting $f(x_{0})$ with $f(x).$ Note that $d_{M}(x_{0},x)=\inf_{\gamma}\|\gamma\|_{M},$ where $\|\gamma\|_{M}$ is the length of $\gamma.$ Then
$d_{N}(f(x_{0}),f(x))$ $\leq$ $\|f\circ\gamma\|_{N}$
$=$ $\int_{0}^{1}\|(f\circ\gamma)^{\prime}\|_{N}(t)dt$
$\leq$ $\int_{0}^{1}\|df_{\gamma(t)}\|\cdot\|\gamma^{\prime}\|_{M}(t)dt,$
where $\|f\circ\gamma\|_{N}$ is the length of $f\circ\gamma.$ Utilizing the conditions, $\{\|df\|\}_{f\in\mathscr{F}}$ is uniformly bounded on $U(x_{0}).$ So, there exists a constant $C>0,$ independent of $x\in U(x_{0})$ and $f\in\mathscr{F},$ such that
$\sup_{f\in\mathscr{F}}\sup_{x\in U(x_{0})}\|df_{x}\|\leq C.$
Therefore, we obtain
$d_{N}(f(x_{0}),f(x))\leq C\int_{0}^{1}\|\gamma^{\prime}\|_{M}(t)dt=C\|\gamma\|_{M}.$
By the arbitrariness of $\gamma,$ it is immediate that
$d_{N}(f(x_{0}),f(x))\leq Cd_{M}(x_{0},x),\ \ \ \ f\in\mathscr{F},$
which deduces that $\mathscr{F}$ is equicontinuous at $x_{0},$ and hence equicontinuous on $M.$ Since $N$ is compact, $\mathscr{F}$ is normal on $M.$
We then prove the necessity. On the contrary, we assume that $\{\|df\|\}_{f\in\mathscr{F}}$ is not locally uniformly bounded on $M.$ Then, there are geodesic balls $B,B^{\prime}$ with $\overline{B}\subset B^{\prime}\subset M,$ and sequences $\{x_{j}\}_{j=1}^{+\infty}\subset\overline{B},$ $\{f_{j}\}_{j=1}^{+\infty}\subset\mathscr{F}$ such that
(8) $\lim_{j\to+\infty}\|(df_{j})_{x_{j}}\|=+\infty.$
Using the compactness of $\overline{B},$ there exists a subsequence $\{x_{j_{k}}\}_{k=1}^{+\infty}\subset\{x_{j}\}_{j=1}^{+\infty}$ which converges to a limit point $x_{0}\in\overline{B}.$ Take a neighborhood $U$ of $x_{0}$ such that $\overline{U}\subset B^{\prime}.$ When $j$ is sufficiently large, we have $x_{j}\in U.$ Moreover, since $\mathscr{F}$ is normal on $M,$ there is a subsequence $\{f_{j_{k}}\}_{k=1}^{+\infty}\subset\{f_{j}\}_{j=1}^{+\infty}$ which converges uniformly to a limit mapping $f_{0}:\overline{U}\to N.$ Invoking Weierstrass theorem, $f_{0}$ is holomorphic on $U$ and we receive
$\lim_{k\to+\infty}\|(df_{j_{k}})_{x}\|=\|(df_{0})_{x}\|\not=+\infty,\ \ \ \ x\in U$
due to the continuity of $\|(df_{0})_{x}\|,\|(df_{j_{k}})_{x}\|$ in $x.$ Since $x_{j_{k}}\to x_{0}$ as $k\to+\infty,$ we conclude that
$\lim_{k\to+\infty}\|(df_{j_{k}})_{x_{j_{k}}}\|=\lim_{k\to+\infty}\|(df_{0})_{x_{j_{k}}}\|=\|(df_{0})_{x_{0}}\|\not$$\infty$.
However, it contradicts with (8). This completes the proof. ∎
Proof of Theorem 1.1
Proof.
Equip $G$ with a left-variant Hermitian metric such that $\{\varepsilon_{1},\cdots,\varepsilon_{m}\}$ forms an orthonormal basis of $\mathfrak{g}.$ Assume that $\mathscr{F}$ is not normal on $\Omega.$ Using Lemma 3.1, for any sequence $\{s_{j}\}_{j=1}^{+\infty}\subset\mathbb{R}$ with $s_{j}\geq 1$ for all $j,$ there exist a compact subset $K\subset\Omega,$ a sequence $\{\tilde{p}_{j}\}_{j=1}^{+\infty}\subset K,$ a sequence $\{\tilde{\xi}_{j}\}_{j=1}^{+\infty}$ with $\tilde{\xi}_{j}\in T^{1,0}_{\tilde{p}_{j}}G$ and $\|\tilde{\xi}_{j}\|_{G}=1,$ as well as a sequence $\{f_{j}\}_{j=1}^{+\infty}\subset\mathscr{F},$ such that
(9) $\big\|(df_{j})_{\tilde{p}_{j}}(\tilde{\xi}_{j})\big\|_{N}\geq j^{s_{j}},\ \ \ \ j=1,2,\cdots$
By the compactness of $K,$ there exists a convergent subsequence of $\{\tilde{p}_{j}\}_{j=1}^{+\infty}.$ Without loss of generality, it is convenient to suppose that $\tilde{p}_{j}\to p_{0}\in K$ as $j\to+\infty$ and $d_{G}(p_{0},\tilde{p}_{j})<1/(2j)$ for all $j,$ in which $d_{G}(\cdot,\cdot)$ is the Riemannian distance on $G.$ Let $B_{G}(p_{0},r)$ be the geodesic ball centered at $p_{0}$ with radius $r$ in $G.$ For $j=1,2,\cdots,$ put
$M_{j}=\max_{g\in\overline{B_{G}(p_{0},1/j)}}\max_{\xi\in T^{1,0}_{g}G,\|\xi\|_{G}=1}\big\|(df_{j})_{g}(\xi)\big\|_{N}.$
Taking $p_{j}\in\overline{B_{G}(p_{0},1/j)}$ and $\xi_{j}\in T^{1,0}_{p_{j}}G$ with $\|\xi_{j}\|_{G}=1$ such that the above expression is maximized, i.e.,
(10) $M_{j}=\big\|(df_{j})_{p_{j}}(\xi_{j})\big\|_{N},\ \ \ \ j=1,2,\cdots$
It is apparent that $p_{j}\to p_{0}$ as $j\to+\infty,$ due to $d_{G}(p_{0},p_{j})\leq 1/j$ inherent in the definition of $M_{j}$. Thus, we assume without loss of generality that $p_{j}\in\Omega$ for all $j.$ Define
$\rho_{j}=\frac{1}{M_{j}},\ \ \ \ j=1,2,\cdots$
It yields from (9) and (10) that
$M_{j}\geq\big\|(df_{j})_{\tilde{p}_{j}}(\tilde{\xi}_{j})\big\|_{N}\geq j^{s_{j}},\ \ \ \ j=1,2,\cdots$
Since $s_{j}\geq 1$ for all $j,$ we arrive at
(11) $0<\rho_{j}\leq j^{-s_{j}}\to 0\ \ \ \ \ \ (j\to+\infty).$
Let us treat the differential of $\exp_{p_{j}}$ at 0:
$(d\exp_{p_{j}})_{\textbf{0}}:\ T^{1,0}_{\textbf{0}}(T^{1,0}_{p_{j}}G)\to T^{1,0}_{p_{j}}G,\ \ \ \ j=1,2,\cdots$
Define
$\phi_{j}(z)=f_{j}\circ\exp_{p_{j}}\big(\alpha_{j}(z)\big),\ \ \ \ j=1,2,\cdots$
where
$\alpha_{j}(z)$ $=$ $\rho_{j}(dL_{p_{j}})_{e}(\varepsilon z^{T}).$
Since $f_{j},\exp_{p_{j}},(dL_{p_{j}})_{e}$ are holomorphic, it is obvious that $\phi_{j}$ is holomorphic. Moreover, since $(dL_{p_{j}})_{e}$ is a linear isometry by (2), we infer that $\alpha_{j}:\mathbb{C}^{m}\to\mathfrak{g}$ is a linear isomorphism with
$\|\alpha_{j}\|=\rho_{j},\ \ \ \ j=1,2,\cdots$
Let $B_{\mathfrak{g}}(r)$ be the geodesic ball centered at 0 with radius $r$ in $\mathfrak{g}.$ Since $\exp_{p_{j}}$ is locally biholomorphic at 0 by Property 2.8, and $(dL_{p_{j}})_{e}$ is linearly isometric, there exists a number $r_{j}>0$ such that
$\exp_{p_{j}}\big((dL_{p_{j}})_{e}\big(\overline{B_{\mathfrak{g}}(r_{j})}\big)\big)\subset B_{G}\big(p_{0},1/j\big).$
It is obvious that $r_{j}\to 0$ as $j\to+\infty.$ Without loss of generality, we assume that $0<r_{j}<1$ for all $j.$ Applying the linear isometry of $(dL_{p_{j}})_{e},$ we have
(12) $\exp_{p_{j}}\big(\alpha(z)\big)\in B_{G}\big(p_{0},1/j\big),\ \ \ \ \|z\|<\frac{r_{j}}{\rho_{j}}.$
Take
$s_{1}=1;\ \ \ \ \ s_{j}=1-\frac{\log r_{j}}{\log j}>1,\ \ \ \ j=2,3,\cdots$
By the identity $x^{1/\log x}=e$ for $x>0,$ we conclude from (11) that $r_{j}\rho^{-1}_{j}\geq j$ for all $j\geq 2.$ By virtue of (12), we deduce that $\phi_{j}$ is well-defined on the ball $\{z\in\mathbb{C}^{m}:\|z\|<j\}$ with $j\geq 2.$ Note that this ball is larger and larger as $j$ increases. Set
$x_{j}(z)=\exp_{p_{j}}(\alpha_{j}(z)),\ \ \ \ j=1,2,\cdots$
As we have noted that $d_{G}(p_{0},p_{j})<1/j$ for all $j.$ Fix a number $r_{0}>0.$ When $\|z\|\leq r_{0}$ and $j$ is large such that $j>r_{0},$ we infer from Theorem 2.9, (2)-(4) and (10) that
$\|(d\phi_{j})_{z}\|$ $=$ $\big\|d\big(f_{j}\circ\exp_{p_{j}}\circ\alpha_{j}\big)_{z}\big\|$
$\leq$ $\big\|(df_{j})_{x_{j}(z)}\big\|\cdot\big\|(d\exp_{p_{j}})_{\alpha_{j}(z)}\big\|\cdot\|(d\alpha_{j})_{z}\|$
$\leq$ $\rho_{j}M_{j}\big\|(d\exp_{p_{j}})_{\alpha_{j}(z)}\big\|\cdot\|z\|$
$\leq$ $r_{0}\frac{e^{C_{\mathfrak{g}}\|\alpha_{j}(z)\|_{G}}-1}{C_{\mathfrak{g}}\|\alpha_{j}(z)\|_{G}}$
$=$ $r_{0}\frac{e^{C_{\mathfrak{g}}\rho_{j}\|z\|}-1}{C_{\mathfrak{g}}\rho_{j}\|z\|}$
$\leq$ $r_{0}\frac{e^{C_{\mathfrak{g}}\rho_{j}r_{0}}-1}{C_{\mathfrak{g}}\rho_{j}r_{0}}$
$\to$ $r_{0}\ \ \ \ \ \ (j\to+\infty).$
To conclude, $\phi_{j}$ is well-defined on larger and larger balls in $\mathbb{C}^{m}$ as $j$ increases, with bounded differential on compact subsets. Using Lemma 3.1, we derive that $\{\phi_{j}\}_{j=1}^{+\infty}$ is normal, i.e., there is a subsequence $\{\phi_{j_{k}}\}_{k=1}^{+\infty}\subset\{\phi_{j}\}_{j=1}^{+\infty}$ such that which converges uniformly on compact subsets of $\mathbb{C}^{m}$ to a holomorphic mapping $\phi:\mathbb{C}^{m}\to N.$ Finally, we remain to examine that $\phi$ is not constant. Since $(d\exp_{p_{j}})_{\textbf{0}}$ is a linear isometry due to Property 2.8, we see that
$(d\exp_{p_{j}})_{\textbf{0}}\circ(d\alpha_{j})_{\textbf{0}}:\ T^{1,0}_{\textbf{0}}\mathbb{C}^{m}\to T^{1,0}_{p_{j}}G,\ \ \ \ j=1,2,\cdots$
are linear isomorphisms. Therefore, there exists a unique vector $t_{j}\in T^{1,0}_{\textbf{0}}\mathbb{C}^{m}$ such that $(d\exp_{p_{j}})_{\textbf{0}}\circ(d\alpha_{j})_{\textbf{0}}(t_{j})=\rho_{j}\xi_{j}$ for all $j.$ By this with (10), we obtain for $j=1,2,\cdots$
$\big\|(d\phi_{j})_{\textbf{0}}(t_{j})\big\|_{N}$ $=$ $\big\|d\big(f_{j}\circ\exp_{p_{j}}(\alpha_{j})\big)_{\textbf{0}}(t_{j})\big\|_{N}$
$=$ $\big\|(df_{j})_{x_{j}(\textbf{0})}\circ(d\exp_{p_{j}})_{\alpha_{j}(\textbf{0})}\circ(d\alpha_{j})_{\textbf{0}}(t_{j})\big\|_{N}$
$=$ $\big\|(df_{j})_{p_{j}}\circ(d\exp_{p_{j}})_{\textbf{0}}\circ(d\alpha_{j})_{\textbf{0}}(t_{j})\big\|_{N}$
$=$ $\rho_{j}\big\|(df_{j})_{p_{j}}(\xi_{j})\big\|_{N}$
$=$ $\rho_{j}M_{j}$
$=$ $1,$
which implies that $\|(d\phi)_{\textbf{0}}\|\not=0.$ Hence, $\phi$ is not a constant mapping.
Assume that the conditions $(a),(b)$ and $(c)$ are satisfied, we show that $\mathscr{F}$ is not normal at $p_{0}.$ Conversely, assume that $\mathscr{F}$ is normal at $p_{0}.$ By Lemma 3.1, there exist $\delta,M>0$ such that $\overline{B_{G}(p_{0},\delta)}\subset\Omega$ and
$\sup_{f\in\mathscr{F}}\|df_{g}\|\leq M,\ \ \ \ g\in\overline{B_{G}(p_{0},\delta)}.$
Fix $z\in\mathbb{C}^{m}.$ Since $p_{j}\to p_{0},$ $0<\rho_{j}\to 0$ as $j\to+\infty,$ when $j$ is large enough, we have $\exp_{p_{j}}(\alpha_{j}(z))\in\overline{B_{G}(p_{0},\delta)}.$ So, it yields from Theorem 2.9 and (2)-(4) that for $j$ large
$\|(d\phi_{j})_{z}\|$ $\leq$ $\|df_{j}\|_{x_{j}(z)}\cdot\big\|(d\exp_{p_{j}})_{\alpha_{j}(z)}\big\|\cdot\|(d\alpha_{j})_{z}\|$
$\leq$ $M\rho_{j}\|z\|\frac{e^{C_{\mathfrak{g}}\|\alpha_{j}(z)\|_{G}}-1}{C_{\mathfrak{g}}\|\alpha_{j}(z)\|_{G}}$
$\leq$ $M\rho_{j}\|z\|\frac{e^{C_{\mathfrak{g}}\rho_{j}\|z\|}-1}{C_{\mathfrak{g}}\rho_{j}\|z\|}$
$\to$ $0\ \ \ \ \ \ (j\to+\infty).$
That is,
$\|d\phi_{z}\|=\lim_{j\to+\infty}\|(d\phi_{j})_{z}\|=0.$
But, the arbitrariness of $z$ implies that $\phi$ is constant, a contradiction. ∎
By the arguments in the proof of Theorem 1.1, it is immediate to deduce that Corollary 1.2 is true.
Proof of Theorem 1.3
Proof.
Equip $G$ with a left-variant Hermitian metric. Conversely, we assume that $\mathscr{F}$ is not normal at some point $p_{0}\in\Omega.$ By Corollary 1.2, there exist sequences $\{f_{j}\}_{j=1}^{+\infty}\subset\mathscr{F},$ $\{p_{j}\}_{j=1}^{+\infty}\subset\Omega$ with $p_{j}\to p_{0}$ as $j\to+\infty,$ $\{\rho_{j}\}_{j=1}^{+\infty}\subset\mathbb{R}$ with $0<\rho_{j}\to 0$ as $j\to+\infty,$ such that $\phi_{j}(z)=f_{j}\circ\exp_{p_{j}}(\theta^{j}z^{T})$ converges uniformly on compact subsets of $\mathbb{C}^{m}$ to a nonconstant holomorphic mapping $\phi:\mathbb{C}^{m}\to N,$ where $\theta^{j}=\rho_{j}(dL_{p_{j}})_{e}(\varepsilon)$ gives a basis of $T^{1,0}_{p_{j}}G$ for each $j.$ Take a number $\delta>0$ such that $\overline{\Delta}\subset\Omega,$ where $\Delta=B_{G}(p_{0},\delta)$ is the geodesic ball centered at $p_{0}$ with radius $\delta$ in $G.$ By $(a),$ we have $(f_{j},\Delta)\in P.$ Further, we get from $(b)$ that $(\phi_{j},D_{j})\in P,$ where $D_{j}=x^{-1}_{j}(\Delta)$ and $x_{j}(z)=\exp_{p_{j}}(\theta^{j}z^{T}).$ Since $0<\rho_{j}\to 0$ as $j\to+\infty,$ $D_{j}$ is larger and larger as $j$ increases, which leads to that $\{\phi_{j};(\phi_{j},D_{j})\in P\}_{j=1}^{+\infty}$ converges uniformly on compact subsets of $\mathbb{C}^{m}$ to $\phi:\mathbb{C}^{m}\to N.$ By $(c),$ we deduce that $(\phi,\mathbb{C}^{m})\in P$ or $(\phi,\mathbb{C}^{m})\in\neg P.$ However, $\phi$ reduces to a constant mapping due to $(d),$ a contradiction. ∎
Finally, we give a Picard theorem of meromorphic functions on connected complex Lie groups.
Theorem 3.2.
Let $G$ be a connected complex Lie group, and $N$ a Brody hyperbolic complex manifold. Then, there exist no nonconstant holomorphic mappings from $G$ into $N.$
Proof.
Since $\exp:\mathfrak{g}\to G$ is locally biholomorphic at $\textbf{0}\in\mathfrak{g},$ we can choose a neighborhood $U$ of 0 in $\mathfrak{g}$ such that the restriction $\exp|_{U}:U\to\exp(U)$ is a biholomorphic mapping. It suffices to verify that any holomorphic mapping $f:G\to N$ is constant on $\exp(U),$ due to the uniqueness theorem of analytic mappings. For an arbitrary $g\in\exp(U),$ there is a unique $X_{g}\in\mathfrak{g}$ such that $\exp(X_{g})=g.$ Treat the following holomorphic curve
$\phi_{g}(z):=f\circ\exp(zX_{g}):\ \mathbb{C}\to N,\ \ \ \ z\in\mathbb{C}.$
Since $N$ is Brody hyperbolic, we have $\phi_{g}\equiv y_{0}$ that is a fixed point in $N.$ So, $f(g)=\phi_{g}(1)=y_{0}.$ On the other hand, with the help of Properties 2.1 and 2.2, we note that $\{\exp(zX_{g})\}_{z\in\mathbb{C}}$ is a holomorphic one-parameter subgroup, which yields immediately that $\phi_{g}(0)=f(e)$ independent of $g.$ Therefore, we receive $y_{0}\equiv f(e)$ for any $g\in\exp(U)$ due to $\phi_{g}(1)=\phi_{g}(0).$ So, $f$ is constant on $\exp(U).$ This completes the proof. ∎
Corollary 3.3.
Let $G$ be a connected complex Lie group. Then, any meromorphic function on $G$ reduces to a constant if it avoids three distinct values.
Corollary 3.4.
There exist no Brody hyperbolic complex Lie groups.
Proof.
Conversely, assume that there exists a Brody hyperbolic complex Lie group $G.$ Let us consider the identity mapping $Id:G\to G,$ which is clearly holomorphic and nonconstant on each connected component of $G.$ However, Theorem 3.2 asserts that $Id$ must be constant on each connected component of $G,$ which is a contradiction. ∎